how do you do this type of problem?
Gary has figured that he needs to take home $30,400 a year to live comfortably. If the government gets 24% of Gary's income, then what must his income be for him to live comfortably?

Answer by stanbon(75887)

Gary has figured that he needs to take home $30,400 a year to live comfortably. If the government gets 24% of Gary's income, then what must his income be for him to live comfortably?
--------------------------
Let his salary be "x"
EQUATION:
x - 0.24x = 30,400
(1-0.24)x = 30,400
0.76x = 30,400
x = $40,000 (Amount of salary he needs)
Cheers,
stan H.